Labor Hours Needed
to Produce

1 Dozen Eggs

1 Pound Ham

Denmark

10

8

Finland

6

4

At which of the following prices, if any, could both Denmark and Finland gain from trade?
Answer #1

Opportunity cost of producing dozen eggs in Denmark = 10/8 = 5/4 pounds of ham.

Opportunity cost of producing dozen eggs in Finland = 6/4 pounds of ham = 3/2 pounds of ham.

When the price lies between 5/4 pounds of ham and 3/2 pounds of ham ,then both Denmark and Finland would gain from trade.

1.4 pounds of ham per dozen eggs lies between this range . Hence,option(C) is correct.